• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Z. Chai

    Log entry time 23:58:18 on June27,2000

    Entry number 45035

    keyword=Swing shift summary


    16:00 Start swing shift. Still having gas problem.
    Mark and Jack are trying to fix the problem.
    The FPP straw efficiency is slowly getting better. But VDC efficiency
    19:35 MCC told us they are having RF problem.
    19:43 Beam back.
    Adjusting the epics bpm offsets to make the epics beam position agree with
    the spot++ result.
    19:58 Beam tripped due to RF problem( South Linac RF problem)
    20:38 MCC told us they are ready to send the beam to Hall A. Ask MCC to adjust
    the BPM Y offsets.
    22:05 Beam back to Hall A (76uA) . With the new bpm offsets for EPICS, the spot++ position agrees with the MCC's set position(from the EPICS).
    From the HANA replay results, the H-Arm VDC efficiency looks fine(as good as before the gas problem), E-Arm VDC efficiency is also better than before.
    22:52 Beam off (RF problem)
    23:29 Beam back.
    ET problems happened several times. Same problem as before.

    Time Run# Events Charge Comment
    16:19 2576 1.22M 167mC
    17:11 2577 0.22M 29mC
    17:23 2578 1.3M 179mC
    18:08 2579 1.3M 179mC
    19:00 2580 1.0M 144mC
    19:49 2581 0.12M 16mC
    22:10 2582 1.07M 146mC FPP straw efficiency back to normal level
    23:29 2583

    /pre>